OpenStudy (anonymous):

If v = <5, -4>, find ||v||. 1 29√ 41√ 21√

geerky42 (geerky42):

\[\Large ||<a,b>|| = \sqrt{a^2+b^2}\]

geerky42 (geerky42):

Can you apply the formula I gave you above here? @davidza6496

OpenStudy (anonymous):

i just plug 5 into A^2 and -4 into B^2

geerky42 (geerky42):

5 into a and -4 into b

geerky42 (geerky42):

so what do you get?

OpenStudy (anonymous):

i got 3?

geerky42 (geerky42):

did you plug in correctly? after plug in, it should looks like this: \[\Large \sqrt{(5)^2+(-4)^2}\]

OpenStudy (anonymous):

yeah i did it again and got 3

OpenStudy (anonymous):

@geerky42

zepdrix (zepdrix):

\[\Large\rm \vec v=\left<\color{orangered}{5},\color{royalblue}{-4}\right>\] \[\Large\rm ||v||=\sqrt{\color{orangered}{5}^2+(\color{royalblue}{-4})^2}\]\[\Large\rm ||v||=\sqrt{25+16}\]Did you remember to square the negative sign?

OpenStudy (anonymous):

ohh ok i did it wrong, but i got it now thank you!

zepdrix (zepdrix):

Cool :)